OpenStudy (anonymous):

What is the value of the expression –c + d when c = –14.9 and d = –2.6? A. –17.5 B. 12.3 C. 17.5 D. –12.3

OpenStudy (anonymous):

B. 12.3 -c + d = -(-14.9) + (-2.6) =14.9 -2.6 =12.3